OpenStudy (anonymous):

what is the cubed root of 121

OpenStudy (anonymous):

Need a calculator for that. 121 is the product of two primes.

hero (hero):

4.946

hero (hero):

approximately

OpenStudy (anonymous):

((11)^2)^1/3 =(11)^2/3 =4.95